OpenStudy (anonymous):

derivative of y=x^(3/2)log(base3)squareroot(x+7)

zepdrix (zepdrix):

\(\large\rm y=x^{3/2}\cdot\log_3\sqrt{x+7}\) This guy? Hmm this looks like a fun one :) lol

OpenStudy (anonymous):

\[y= x ^{3/2}\log_{3}\sqrt{x+7}\]

OpenStudy (anonymous):

yeah lol

zepdrix (zepdrix):

Do you remember your change of base formula for logs?\[\large\rm \log_{a}(b)=\frac{\ln(b)}{\ln(a)}\]We would choose natural log like this ^ Because it's easier to differentiate.

zepdrix (zepdrix):

Unless you have a special formula memorized for log derivatives :) \[\large\rm y=x^{3/2}\cdot\frac{\ln\sqrt{x+7}}{\ln3}\]

OpenStudy (anonymous):

so it would be \[(3/2)x ^{1/2}*(\ln \sqrt{x+7}/\ln3)\]

OpenStudy (anonymous):

?

zepdrix (zepdrix):

No no no :O Product rule.

OpenStudy (anonymous):

what should i do then?

zepdrix (zepdrix):

I'll do the "set up" and you can let me know if it makes sense:\[\large\rm y'=\color{royalblue}{\left(x^{3/2}\right)'}\cdot\frac{\ln\sqrt{x+7}}{\ln3}+x^{3/2}\cdot\color{royalblue}{\left(\frac{\ln\sqrt{x+7}}{\ln3}\right)'}\]This is our product rule "set up". I haven't taken any derivatives yet. I'm just setting it up. We need to differentiate the blue parts.

OpenStudy (anonymous):

is the first blue part \[(3/2)x ^{(3/2)-1}\]

zepdrix (zepdrix):

\[\large\rm y'=\color{orangered}{\left(\frac{3}{2}x^{1/2}\right)}\cdot\frac{\ln\sqrt{x+7}}{\ln3}+x^{3/2}\cdot\color{royalblue}{\left(\frac{\ln\sqrt{x+7}}{\ln3}\right)'}\]Ok ya that sounds right!

zepdrix (zepdrix):

This other guy is a bit tricky though, ya? :)

OpenStudy (anonymous):

\[1/((1/2)x ^{-1/2}+(1/2)7^{-1/2}) /(1/3)\]

OpenStudy (anonymous):

?

zepdrix (zepdrix):

No. And oh boy, did you ever make things complicated :D lol

zepdrix (zepdrix):

Let's back up a sec, apply a helpful log rule before we take the derivative.

OpenStudy (anonymous):

sorry lol

OpenStudy (anonymous):

\[\ln \sqrt{x+7}-\ln 3\]

OpenStudy (anonymous):

?

zepdrix (zepdrix):

First notice that ln3 is just a constant, so we can pull it out of the differentiation process and ignore it.\[\large\rm \color{royalblue}{\left(\frac{\ln\sqrt{x+7}}{\ln3}\right)'}=\frac{1}{\ln3}\color{royalblue}{\left(\ln\sqrt{x+7}\right)'}\]

OpenStudy (anonymous):

what about the x+7

zepdrix (zepdrix):

Next rewrite your square root as a 1/2 power,\[\large\rm \frac{1}{\ln3}\color{royalblue}{\left(\ln(x+7)^{1/2}\right)'}\]And apply another log rule,\[\large\rm =\frac{1}{\ln3}\color{royalblue}{\left(\frac{1}{2}\ln(x+7)\right)'}\]And again pull the 1/2 out front because it's just a constant,\[\large\rm =\frac{1}{2\ln3}\color{royalblue}{\left(\ln(x+7)\right)'}\]The problem becomes a lot easier to deal with at this point.

zepdrix (zepdrix):

This is one of the things that can make Calc a difficult subject, remember all of these silly Algebra rules :) They're really helpful though.

OpenStudy (anonymous):

does the x+7 become 8

OpenStudy (anonymous):

or does it become 2

zepdrix (zepdrix):

I don't understand what you're asking

OpenStudy (anonymous):

does \[(\ln(x+7))\prime = \ln(2)\prime = 1/2\]

zepdrix (zepdrix):

?? 0_o

zepdrix (zepdrix):

I have no idea where that 2 is coming from.. this is so confusing

zepdrix (zepdrix):

Do you remember your derivative for ln(x) ?

OpenStudy (anonymous):

isnt the derivative of a constant = 1?

OpenStudy (anonymous):

so the derivative of x = 1 and the derivative of 7 = 1

OpenStudy (anonymous):

so 1+1 = 2

OpenStudy (anonymous):

???

zepdrix (zepdrix):

The derivative of a constant is = 0. x is not a constant. But let's not worry about that. That's not what we're dealing with here. I'm asking if you know the derivative of \(\large\rm \ln(x)\)

OpenStudy (anonymous):

is it 1?

zepdrix (zepdrix):

... No :(

OpenStudy (anonymous):

isnt the derivative of ln(x) = 1/x

zepdrix (zepdrix):

Yes.

zepdrix (zepdrix):

So then by that same rule,\[\large\rm \frac{d}{dx}\ln(x+7)=\frac{1}{(x+7)}\]

OpenStudy (anonymous):

wait so is the answer \[(3/2)x ^{1/2}-((\ln(\sqrt{x+7})/\ln3)*x ^{3/2})*(1/x+7)\]

zepdrix (zepdrix):

No :(

OpenStudy (anonymous):

what is the right answer?

zepdrix (zepdrix):

By the steps I was showing you before, the derivative of that second blue thing we were working on, would be,\[\large\rm \color{royalblue}{\left(\frac{\ln\sqrt{x+7}}{\ln3}\right)'}=\color{orangered}{\frac{1}{2\ln3(x+7)}}\]

zepdrix (zepdrix):

\[\large\rm y'=\color{orangered}{\left(\frac{3}{2}x^{1/2}\right)}\cdot\frac{\ln\sqrt{x+7}}{\ln3}+x^{3/2}\cdot\color{royalblue}{\left(\frac{\ln\sqrt{x+7}}{\ln3}\right)'}\] So I guess our final result is something like this,\[\large\rm y'=\color{orangered}{\left(\frac{3}{2}x^{1/2}\right)}\cdot\frac{\ln\sqrt{x+7}}{\ln3}+x^{3/2}\cdot\color{orangered}{\frac{1}{(2\ln3)(x+7)}}\]

zepdrix (zepdrix):

Practice more problems cp :O Really really need to study!

OpenStudy (anonymous):

OH OKAY

OpenStudy (anonymous):

I GET IT NOW

OpenStudy (anonymous):

THANK YOU SO MUCH